OpenStudy (anonymous):

Hi All, I need help on another equation. I know the answer but it's the process once again I can't get: 6-3x=5x-10x+18. Any help would be greatly appreciated, Denise

OpenStudy (anonymous):

Oh hi Luis:)

OpenStudy (anonymous):

How did you get that Luis, the 2x first of all? did you combine like terms?

OpenStudy (anonymous):

so 6 + 18 first right?

OpenStudy (anonymous):

then -10x + -3x?

OpenStudy (anonymous):

Hi Radar, you always seem to come to my rescue, LOL!!

OpenStudy (radar):

6-3x=5x-10x+ 18 combine similar terms on each side of the equal sign 6-3x=-5x+18 add 5x to both sides. 6+2x=18 subtract 6 from both sides getting 2x=12 divide both sides by 2 x=6

OpenStudy (anonymous):

Ok, I am reading it, I'll get it but I won't quit my day-job:(

OpenStudy (anonymous):

omgosh, it was so easy, I just wish I could retain this stuff.

OpenStudy (radar):

Hi nwcountry, to isolate x, you have to do it algebraically by "legally" manipulating the equation getting the unknown on one side, and the numerical values on the other. You do this by either adding, subtracting, multiplying or dividing both sides of the equal sign until you end up with "x" on one side and the value on the other. Remember, while doing this, you must do the same to both sides or otherwise you will unbalance the equation and get a wrong answer.

OpenStudy (anonymous):

I understand, ok radar. We have worked these in class and in homework but I forget things and then there is a test tomorrow, a final and I am reviewing everything we've done in a very, short 9 weeks:( That's life hey. One more question, do you start isolating on the right side of the equal every time on these, like you started with adding the 5x to both sides first?

OpenStudy (radar):

Personally I try and get the "literal variable" (the x or y or whatever) on the left, and the numerical values on the right of the equal sign. The equal sign is like the balance point it is saying what is on the left is the same value as what is on the right. It is balanced like the old time balance scales. If you add on the left, you must do the same to the right or you will upset the balance.

OpenStudy (radar):

I should point, you could just as easily get the x or y isolated to the right and the numerical value to the left.

OpenStudy (anonymous):

Ok, I do remember that in class as well! Thank you for all your help. I think practice is the ticket. Thanks again and I will ttyl, Denise PS yes, but I too like the X or variable on the left side:) Maybe see you tomorrow although after my final, I don't know that I will want to do Math, LOL!! I love math but it is getting harder every term. Nite Radar;)

OpenStudy (radar):

When I added the 5x to both sides, I was starting my effort to get the x isolated to the left side.

OpenStudy (anonymous):

Yes, you are very good at this stuff;)

OpenStudy (radar):

Good night Denise, I do know it is earlier in Oregon than in MO. lol

OpenStudy (anonymous):

You do it just like my instructor and I think he is very good. He even gets frustrated that we have such a short time to learn such a lot.

OpenStudy (anonymous):

Yes, you are up late;)

OpenStudy (radar):

Luis is good, but he is brief. And as you say "practice" that is the ticket.

OpenStudy (anonymous):

I will do my best:) Yes, Luis is good as well. I like the site, so many helpful people. I will be glad if I can help someone else one day:)

OpenStudy (radar):

Good luck with your studies, I'm sure you'll do O.K. on the test.

OpenStudy (anonymous):

I so pray I do, it really is important I do well on this one:) I better go but really enjoyed talking to you and I appreciate how you take the time to show me all the steps, Denise:)

OpenStudy (anonymous):

nite radar:)